OpenStudy (anonymous):

Use common logarithms or natural logarithms and a calculator to evaluate to four decimal places. log4 20

OpenStudy (paxpolaris):

\[\Large \log_a(x) = {\log_b(x) \over \log_b(a) }\]

OpenStudy (paxpolaris):

so, \[\large\log_4(20) = {\log 20 \over \log 4}\]OR\[\large\log_4(20) = {\ln 20 \over \ln 4}\]